# Multitask-based Evaluation of Open-Source LLM on Software Vulnerability

### v2

@@ Abstract
This paper proposes a pipeline for quantitatively evaluating interactive LLMs using publicly available datasets. We carry out an extensive technical evaluation of LLMs using Big-Vul covering four different common software vulnerability tasks. We evaluate the multitask and multilingual aspects of LLMs based on this dataset. We find that the existing state-of-the-art methods are generally superior to LLMs in software vulnerability detection. Although LLMs improve accuracy when providing context information, they still have limitations in accurately predicting severity ratings for certain CWE types. In addition, LLMs demonstrate some ability to locate vulnerabilities for certain CWE types, but their performance varies among different CWE types. Finally, LLMs show uneven performance in generating CVE descriptions for various CWE types, with limited accuracy in a few-shot setting. Overall, though LLMs perform well in some aspects, they still need improvement in understanding the subtle differences in code vulnerabilities and the ability to describe vulnerabilities to fully realize their potential. Our evaluation pipeline provides valuable insights for further enhancing LLMs’ software vulnerability handling capabilities.
@@ 1 Introduction
Recently, Large Language Models (LLMs) [1] have been widely adopted since the advances in Natural Language Processing (NLP) which enable LLM to be well-trained with both billions of parameters and billions of training samples, consequently bringing a large performance improvement on tasks adopted by LLMs. LLMs can be easily used for a downstream task by being fine-tuned [2] or being prompted [3] since they are trained to be general and they can capture different knowledge from various domain data. Fine-tuning is used to update model parameters for a particular downstream task by iterating the model on a specific dataset while prompting can be directly used by providing natural language descriptions or a few examples of the downstream task. Compared to prompting, fine-tuning is expensive since it requires additional model training and has limited usage scenarios, especially in cases where sufficient training datasets are unavailable. LLMs have demonstrated remarkable language comprehension and generation capabilities, and have been able to perform well on a variety of natural language processing tasks, such as text summarization [4]. Given the outstanding performance of LLMs, there is a growing focus on exploring their potential in software engineering tasks and seeking new opportunities to address them. Currently, as more and more LLMs designed for software engineering tasks are deployed [5, 6, 7, 8, 9, 10, 11], many research works focused on the application of LLMs in the software engineering domain [12, 13, 14, 15, 16]. However, in the existing literature, adequate systematic reviews and surveys have been conducted on LLMs in areas such as generating high-quality code and high-coverage test cases [17, 18], but a systematic review and evaluation of LLMs in the field of software vulnerability is still missing.
@@
  • RQ-1: How does LLMs perform on vulnerability detection? Software Vulnerabilities (SVs) can expose software systems to risk situations and consequently software function failure. Therefore, detecting these SVs is an important task for software quality assurance. We want to explore the ability of LLMs on vulnerability detection as well as the performance difference compared with state-of-the-art approaches.
  • RQ-2: How does LLMs perform on vulnerability assessment? In practice, due to the limitation of SQA resources [19], it is impossible to treat all detected SVs equally and fix all SVs simultaneously. Thus, it is necessary to prioritize these detected software vulnerabilities for better treatment. An effective solution to prioritize those SVs is to use one of the most widely known SV assessment frameworks CVSS (Common Vulnerability Scoring System) [20], which characterizes SVs by considering three metric groups: Base, Temporal, and Environmental. The metrics that are in the groups can be further used as the criterion for selecting serious SVs to fix early. Therefore, we want to explore the ability of LLMs to assess vulnerabilities.
  • RQ-3: How do LLMs perform on vulnerability location? Identifying the precise location of vulnerabilities in software systems is of critical importance for mitigating risks and improving software quality. The vulnerability location task involves pinpointing these weaknesses accurately and helps to narrow the scope for developers to fix problems. Therefore, we aim to investigate LLMs’ capability in effectively identifying the precise location of vulnerabilities in software systems.
  • RQ-4: How does LLMs perform on vulnerability description? Understanding the intricacies of vulnerabilities in software systems plays a pivotal role in alleviating risks and bolstering software quality. The vulnerability description task focuses on conveying a detailed explanation of these identified issues in the source codes and helps participants to better understand the risk as well as its impacts. We aim to assess LLMs’ capacity to effectively generate the description of vulnerabilities within software systems.
@@
TABLE I: Insights and Takeaways: Evaluating LLMs on Software Vulnerability
Dimension Findings or Insights
Vulnerability Detection 1 . Fine-tuned LLMs perform weaker than transformer-based methods, yet comparably to graph-based methods. Moreover, LLMs in the few-shot setting show lower performance than existing methods. 2. After fine-tuning, the detection capability of LLMs has improved, except for Mistral. Larger models usually perform better, but performance can also be influenced by model design and pre-training data. 3. WizardCoder has the best vulnerability detection capability, while Mistral is the worst.
Vulnerability Assessment 4 . Larger model parameter counts did not enhance vulnerability assessment performance with LLMs, prioritize smaller parameter models for better cost-performance balance. 5. LLMs have a limited capacity for assessment of vulnerability severity based on source code only, but can be extremely improved if provided with more context information in most cases.
Vulnerability Localization 6 . Few-shot setting expose LLM limitations, but fine-tuning enhances caution. 7. Mistral’s significant improvement after fine-tuning showcases its potential.
Vulnerability Description 8 . CodeLlama, StarCoder, WizardCoder, and Mistral excel at learning from historical description data.
@@ 2.2 Software Vulnerability
Software Vulnerabilities (SVs) can expose software systems to risk situations and consequently make the software under cyber-attacks, eventually causing huge economic losses and even threatening people’s lives. Therefore, vulnerability databases have been created to document and analyze publicly known security vulnerabilities. For example, Common Vulnerabilities and Exposures (CVE) [33, 34] and SecurityFocus [35] are two well-known vulnerability databases. Besides, Common Weakness Enumeration (CWE) defines the common software weaknesses of individual vulnerabilities, which are often referred to as vulnerability types of CVEs. To better address these vulnerabilities, researchers have proposed many approaches for understanding the effects of software vulnerabilities, including SV detection [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], SV assessment [47, 48, 20, 49, 50], SV localization [51, 52, 53], SV repair [54, 55, 56, 57] as well as SV description [58, 59, 60, 61]. Many novel technologies are adopted to promote the progress of software vulnerability management, including software analysis [62, 63], machine learning [36, 43], and deep learning [47, 52], especially large language models [59, 60].
@@ 3 Experimental Design
In this section, we present our studied dataset, our studied LLMs, the prompt engineering, the baseline approaches, the evaluation metrics, and the experiment settings.
@@ 3.1 Studied Dataset
In our work, some baselines need to obtain the structure information (e.g., control flow graph (CFG), data flow graph (DFG)) of the studied functions. Therefore, we adopt the same toolkit with Joern [64] to transform functions. The functions are dropped out directly if they cannot be transformed by Joern successfully. We also remove the duplicated functions and the statistics of the studied dataset are shown in Table II. Finally, the filtered dataset is used for evaluation. We follow the same strategy to build the training data, validating data, and testing data from the original dataset with previous work does [37, 54, 46]. Specifically, 80% of functions are treated as training data, 10% of functions are treated as validation data, and the left 10% of functions are treated as testing data. We also keep the distribution as same as the original ones in training, validating, and testing data.
TABLE II: The statistic of studied dataset
@@
Training 4,208 4,208 8,416 1
Validating 526 9,631 10,157 0.055
Testing 526 9,631 10,157 0.055
@@ 3.2 Studied LLMs
TABLE III: Overview of studied LLMs
@@
Group 1: Code-related LLMs.
DeepSeek-Coder developed by DeepSeek AI [9] is composed of a series of code language models, each trained from scratch on 2T tokens, with a composition of 87% code and 13% natural language in both English and Chinese. They provide various sizes of the code model, ranging from 1B to 33B versions. Each model is pre-trained on project-level code corpus by employing a window size of 16K and an extra fill-in-the-blank task, to support project-level code completion and infilling. For coding capabilities, DeepSeek-Coder achieves state-of-the-art performance among open-source code models on multiple programming languages and various benchmarks.
@@
Phi-2 proposed by Microsoft [67] packed with 2.7 billion parameters. It’s designed to make machines think more like humans and do it safely. Phi-2 isn’t just about numbers; it’s about a smarter, safer way for computers to understand and interact with the world. Phi-2 stands out because it’s been taught with a mix of new language data and careful checks to make sure it acts right. It’s built to do many things like writing, summarizing texts, and coding, but with better common sense and understanding than its earlier version, Phi-1.5. Phi-2’s evaluation demonstrates its proficiency over larger models in aggregated benchmarks, emphasizing the potential of smaller models to achieve comparable or superior performance to their larger counterparts. This is particularly evident in its comparison with Google Gemini Nano 2, where Phi-2 outshines despite its smaller size.
3.3 Prompt Engineering
We follow the prompt similar to those used in the artifacts, papers, or technical reports associated with each corresponding model [5, 10, 11]. For fine-tuning setting and few-shot setting, we use the same prompt, where each prompt contains three pieces of information: (1) task description, (2) source code, and (3) indicator. Using the software vulnerability detection task as an example, the prompt utilized for LLM consists of three crucial components, as depicted in Fig. 3:
  • Task Description (marked as ①). LLM is provided with the description constructed as ‘‘If this C code snippet has vulnerabilities, output Yes; otherwise, output No’’. The task descriptions used in the SV detection task vary based on the source programming language we employ.
  • Indicator (marked as ③). LLM is instructed to think about the results. In this paper, we follow the best practice in previous work [12] and adopt the same prompt named ‘‘// Detection’’.
@@ 3.4 Baselines
To comprehensively compare the vulnerability detection performance of LLMs with existing state-of-the-art (SOTA) approaches, in this study, we consider the five approaches: Devign [36], ReVeal [45], IVDetect [52], LineVul [37], and SVulD [46]. We briefly introduce them as follows.
@@ 3.5 Evaluation Metrics
For considered software vulnerability-related tasks, we will perform evaluations using the widely adopted performance metrics. More precisely, to evaluate the effectiveness of LLMs on vulnerability detection and vulnerability assessment, we consider the following four metrics: F1-score, Recall, Precision, and Accuracy.
For the software vulnerability location task, we adopt the Hit@Acc, Precision, and Recall metrics (refer to Section 4.3 for more details). For the software vulnerability description task, we use Rouge-1, Rouge-2, and Rouge-L metrics.

### v3

@@ Abstract
This paper proposes a pipeline for quantitatively evaluating interactive Large Language Models (LLMs) using publicly available datasets. We carry out an extensive technical evaluation of LLMs using Big-Vul covering four different common software vulnerability tasks. This evaluation assesses the multi-tasking capabilities of LLMs based on this dataset. We find that the existing state-of-the-art approaches and pre-trained Language Models (LMs) are generally superior to LLMs in software vulnerability detection. However, in software vulnerability assessment and location, certain LLMs (e.g., CodeLlama and WizardCoder) have demonstrated superior performance compared to pre-trained LMs, and providing more contextual information can enhance the vulnerability assessment capabilities of LLMs. Moreover, LLMs exhibit strong vulnerability description capabilities, but their tendency to produce excessive output significantly weakens their performance compared to pre-trained LMs. Overall, though LLMs perform well in some aspects, they still need improvement in understanding the subtle differences in code vulnerabilities and the ability to describe vulnerabilities to fully realize their potential. Our evaluation pipeline provides valuable insights into the capabilities of LLMs in handling software vulnerabilities.
@@ 1 Introduction
Recently, Large Language Models (LLMs) [1] have been widely adopted since the advances in Natural Language Processing (NLP) which enable LLM to be well-trained with both billions of parameters and billions of training samples, consequently bringing a large performance improvement on tasks adopted by LLMs. LLMs can be easily used for a downstream task by being fine-tuned [2] or being prompted [3] since they are trained to be general and they can capture different knowledge from various domain data. Fine-tuning is used to update model parameters for a particular downstream task by iterating the model on a specific dataset while prompting can be directly used by providing natural language descriptions or a few examples of the downstream task. Compared to prompting, fine-tuning is expensive since it requires additional model training and has limited usage scenarios, especially in cases where sufficient training datasets are unavailable.
LLMs have demonstrated remarkable language comprehension and generation capabilities, and have been able to perform well on a variety of natural language processing tasks, such as text summarization [4]. Given the outstanding performance of LLMs, there is a growing focus on exploring their potential in software engineering tasks and seeking new opportunities to address them. Currently, as more and more LLMs designed for software engineering tasks are deployed [5, 6, 7, 8, 9, 10, 11], many research works focused on the application of LLMs in the software engineering domain [12, 13, 14, 15, 16]. However, in the existing literature, adequate systematic reviews and surveys have been conducted on LLMs in areas such as generating high-quality code and high-coverage test cases [17, 18], but a systematic review and evaluation of open-source LLMs in the field of software vulnerability is still missing.
@@
  • RQ-1: How do LLMs perform on vulnerability detection? Software Vulnerabilities (SVs) can expose software systems to risk situations and consequently software function failure. Therefore, detecting these SVs is an important task for software quality assurance. We aim to explore the ability of LLMs on vulnerability detection as well as the performance difference compared with state-of-the-art approaches and pre-trained Language Models (LMs).
  • RQ-2: How do LLMs perform on vulnerability assessment? In practice, due to the limitation of SQA resources [19], it is impossible to treat all detected SVs equally and fix all SVs simultaneously. Thus, it is necessary to prioritize these detected software vulnerabilities for better treatment. An effective solution to prioritize those SVs is to use one of the most widely known SV assessment frameworks CVSS (Common Vulnerability Scoring System) [20], which characterizes SVs by considering three metric groups: Base, Temporal, and Environmental. The metrics that are in the groups can be further used as the criterion for selecting serious SVs to fix early. Therefore, we aim to explore the ability of LLMs to assess vulnerabilities and compare their performance with pre-trained LMs.
  • RQ-3: How do LLMs perform on vulnerability location? Identifying the precise location of vulnerabilities in software systems is of critical importance for mitigating risks and improving software quality. The vulnerability location task involves pinpointing these weaknesses accurately and helps to narrow the scope for developers to fix problems. Therefore, we aim to investigate LLMs’ capability in effectively identifying the precise location of vulnerabilities in software systems, alongside evaluating their performance against state-of-the-art approaches and pre-trained LMs.
  • RQ-4: How do LLMs perform on vulnerability description? The vulnerability description task focuses on conveying a detailed explanation of these identified issues in the source codes and helps participants to better understand the risk as well as its impacts. Understanding the intricacies of vulnerabilities in software systems plays a pivotal role in alleviating risks and bolstering software quality. The vulnerability description task focuses on conveying a detailed explanation of these identified issues in the source codes and helps participants to better understand the risk as well as its impacts. Our goal is to evaluate LLMs’ ability to effectively generate vulnerability descriptions within software systems and compare their performance with that of pre-trained LMs.
@@
TABLE I: Takeaways: Evaluating LLMs on Software Vulnerability
Dimension Finding or Guidance
Vulnerability Detection 1 . LLMs can detect vulnerabilities, but fine-tuned LLMs perform weaker than transformer-based approaches. Considering the computational resources and time costs of deploying LLMs, transformer-based approaches for vulnerability detection are a more efficient choice. 2. After fine-tuning, the detection capability of LLMs has improved. Larger models usually perform better, but performance can also be influenced by model design and pre-training data. Therefore, fine-tuning the LLM on domain-specific data before using it as a vulnerability detector is necessary. 3. In general, different LLMs complementing each other, while CodeLlama obtains better performance in terms of F1-score, Precision, and Recall.
Vulnerability Assessment 4 . Overall, fine-tuned code-related LLMs outperform pre-trained language models in vulnerability assessment. When resources permit, fine-tuning DeepSeek-Coder 6.7B for vulnerability assessment is optimal, as it outperforms the pre-trained language models across four metrics. 5. LLMs have the capacity for assessment of vulnerability severity based on source code, and can be improved by providing more context information.
Vulnerability Location 6 . Few-shot setting exposes LLM’s limitations, and fine-tuning can greatly enhance the vulnerability location capabilities of LLMs. 7. Fine-tuning code-related LLMs as vulnerability locators is beneficial, as they can outperform pre-trained language models in terms of F1-score, precision, and FPR.
Vulnerability Description 8 . LLMs exhibit significantly weaker performance in generating vulnerability descriptions compared to pre-trained language models. Therefore, fine-tuning pre-trained language models for vulnerability detection is recommended.
@@ 2.1 Large Language Model
The application of LLMs in software engineering has seen a surge, with models like ChatGPT being employed for various tasks (e.g., code review, code generation, and vulnerability detection). Although some works use LLMs for vulnerability tasks [33, 34], our work differs from these previous studies in the following aspects. (1) Closed-source ChatGPT vs. Open-source LLMs: They only explore the capabilities of the closed-source ChatGPT in vulnerability tasks, whereas we investigate the abilities of both open-source code-related LLMs and general LLMs in these tasks. (2) Prompts vs. Few-shot and Fine-tuning Settings: They focus solely on the performance of LLMs using prompts, which introduces randomness and hinders the reproducibility of their findings. In contrast, we examine the capabilities of LLMs under both few-shot and fine-tuning settings, providing the source code and corresponding model files to ensure the reproducibility of our experimental results.
@@ 2.2 Software Vulnerability
Software Vulnerabilities (SVs) can expose software systems to risk situations and consequently make the software under cyber-attacks, eventually causing huge economic losses and even threatening people’s lives. Therefore, vulnerability databases have been created to document and analyze publicly known security vulnerabilities. For example, Common Vulnerabilities and Exposures (CVE) [35, 36] and SecurityFocus [37] are two well-known vulnerability databases. Besides, Common Weakness Enumeration (CWE) defines the common software weaknesses of individual vulnerabilities, which are often referred to as vulnerability types of CVEs. To better address these vulnerabilities, researchers have proposed many approaches for understanding the effects of software vulnerabilities, including SV detection [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], SV assessment [51, 52, 20, 53, 54], SV location [55, 56, 57], SV repair [58, 59, 60, 61] as well as SV description [62, 63, 64, 65]. Many novel technologies are adopted to promote the progress of software vulnerability management, including software analysis [66, 67], machine learning [38, 45], and deep learning [51, 56], especially LLMs [63, 64].
@@ 3 Experimental Design
In this section, we present our studied dataset, our studied LLMs, the techniques for fine-tuning, the prompt engineering, the baseline approaches, the evaluation metrics, and the experiment settings.
@@ 3.1 Studied Dataset
We follow the same strategy to build the training data, validation data, and testing data from the original dataset with previous work does [39, 58, 48]. Specifically, 80% of functions are treated as training data, 10% of functions are treated as validation data, and the left 10% of functions are treated as testing data. We also keep the distribution as same as the original ones in training, validation, and testing data. Notice that we undersample the non-vulnerable functions to produce approximately balanced training data at the function level, while the validation and testing data remain in the original imbalanced ratio. To clean and normalize the dataset, we remove empty lines, leading and trailing spaces in each line, as well as comments from the source code. Finally, the split dataset is used for evaluation and the statistics are shown in Table II.
TABLE II: Statistic of the studied dataset
@@
Training 8,720 8,720 17,440 1
Validation 1,090 17,774 18,864 0.061
Testing 1,090 17,774 18,864 0.061
We undersample the non-vulnerable functions to produce approximately balanced training data.
@@ 3.2 Studied LLMs
TABLE III: Overview of the studied LLMs
@@
Group 1: Code-related LLMs. DeepSeek-Coder developed by DeepSeek AI [9] is composed of a series of code language models, each trained from scratch on 2T tokens, with a composition of 87% code and 13% natural language in both English and Chinese. They provide various sizes of the code model, ranging from 1B to 33B versions. Each model is pre-trained on project-level code corpus by employing a window size of 16K and an extra fill-in-the-blank task, to support project-level code completion and infilling. For coding capabilities, DeepSeek-Coder achieves state-of-the-art performance among open-source code models on multiple programming languages and various benchmarks.
@@
Phi-2 proposed by Microsoft [70] packed with 2.7 billion parameters. It is designed to make machines think more like humans and do it safely. Phi-2 is not just about numbers; it is about a smarter, safer way for computers to understand and interact with the world. Phi-2 stands out because it is been taught with a mix of new language data and careful checks to make sure it acts right. It is built to do many things like writing, summarizing texts, and coding, but with better common sense and understanding than its earlier version, Phi-1.5. Phi-2’s evaluation demonstrates its proficiency over larger models in aggregated benchmarks, emphasizing the potential of smaller models to achieve comparable or superior performance to their larger counterparts. This is particularly evident in its comparison with Google Gemini Nano 2, where Phi-2 outshines despite its smaller size.
3.3 Model Fine-Tuning
The four software vulnerability tasks can be categorized into two types: discriminative task (i.e., software vulnerability detection, software vulnerability assessment, and software vulnerability location) and generative task (i.e., software vulnerability description). Therefore, fine-tuning LLMs for software vulnerability tasks can be undertaken through both discriminative and generative methods, each method specifically designed to make LLMs aligned with the task. In particular, we treat the discriminative tasks as binary classification, while treating the generative task as generation one. The architectures for the two paradigms are presented in Fig. LABEL:fig:overview.
Discriminative Fine-Tuning. For vulnerability detection and vulnerability assessment, we utilize the “AutoModelForSequenceClassification” class provided by the Transformers library to implement discriminative fine-tuning. “AutoModelForSequenceClassification” is a generic model class that will be instantiated as one of the sequence classification model classes of the library when created with the “AutoModelForSequenceClassification.from_pretrained(model_name_or_path)” class method.
For vulnerability location, we follow previous works [72, 73] that use LLMs to classify individual code lines as either vulnerable or non-vulnerable. For a token sequence $T=\{t_{1},t_{2},...,t_{n}\}$ of the function, the model’s decoder component, denoted as $M$, processes $T$ to yield a sequence of output vectors: $O=M(T)=\{o_{1},o_{2},...,o_{L}\}$, where $O$ represents the output tensor with dimensions $L\times H$, $L$ signifies the sequence length, and $H$ denotes the hidden dimension size. During the process, the contextual information is captured by the masked self-attention mechanisms in the decoder of LLMs, where masked self-attention limits the sight to the preceding part of tokens. Each output vector $o_{i}$ that represents the last token of one line is subsequently associated with a label (i.e., 0 or 1). The optimization process employs the binary cross-entropy as the loss function.
@@ 3.3 Model Fine-Tuning
Generative Fine-Tuning. Generative fine-tuning aims to equip LLMs with the ability to perform Sequence-to-Sequence (Seq2Seq) tasks. Specifically, this involves inputting vulnerable code and generating the corresponding CVE descriptions related to the vulnerabilities. To calculate the loss during fine-tuning, we utilize the cross-entropy loss function, which is commonly used for Seq2Seq tasks. In this context, the loss measures the difference between the generated output sequence and the target sequence.
3.4 Prompt Engineering
For few-shot setting, we follow the prompt similar to those used in the artifacts, papers, or technical reports associated with each corresponding model [5, 10, 11], where each prompt contains three pieces of information: (1) task description, (2) source code, and (3) indicator. Using the software vulnerability detection task as an example, the prompt utilized for LLM consists of three crucial components, as depicted in Fig. 5:
  • Task Description (marked as ①). We provide LLM with the description constructed as ‘‘If this C code snippet has vulnerabilities, output Yes; otherwise, output No’’. The task description used in the SV detection task varies based on the source programming language we employ.
  • Indicator (marked as ③). We instruct LLM to think about the results. In this paper, we follow the best practice in previous work [12] and adopt the same prompt named ‘‘// Detection’’.
@@ 3.5 Baselines
To comprehensively compare the performance of LLMs with existing approaches, in this study, we consider the various pre-trained Language Models (LMs). As shown in Table V, these models have fewer than 220 million parameters and can be categorized into two categories: encoder-only LMs and encoder-decoder LMs. Encoder-only LMs (i.e., CodeBERT [24], GraphCodeBERT [25], and UniXcoder [26]) contain only the encoder component of a Transformer. They are designed for learning data representations and trained using the Masked Language Modeling (MLM) objective. Encoder-decoder LMs (i.e., PLBART [27], CodeT5 [7], and CodeT5+ [8]) have been proposed for sequence-to-sequence tasks. They are trained to recover the correct output sequence given the original input, often through span prediction tasks where random spans are replaced with artificial tokens. Recently, researchers have combined MLM with generative models for bidirectional and autoregressive text generation or infilling [74]. All these LMs can potentially be used for our tasks, so we evaluate these LMs.
TABLE V: Overview of the studied LMs
Models # Para. Model Type Models # Para. Model Type
CodeBERT 125M Encoder-only LM PLBART 140M Encoder-decoder LM
GraphCodeBERT 125M Encoder-only LM CodeT5 220M Encoder-decoder LM
UniXcoder 125M Encoder-only LM CodeT5+ 220M Encoder-decoder LM
For UniXcoder, we use encoder-only mode.
For vulnerability location, we also consider Devign [38], Reveal [47], IVDetect [56], and LineVul [39] as baselines. In addressing vulnerability detection, we also include SVulD [48] in addition to the aforementioned approaches. We briefly introduce them as follows.
@@ 3.6 Evaluation Metrics
For considered software vulnerability-related tasks, we will perform evaluations using the widely adopted performance metrics. More precisely, to evaluate the effectiveness of LLMs on vulnerability detection and vulnerability assessment, we consider the following four metrics: F1-score, Recall, Precision, and Accuracy. Additionally, for vulnerability location, besides the four aforementioned metrics, we also consider the Top-k Accuracy and FPR metrics. For vulnerability description, we use Rouge-1, Rouge-2, and Rouge-L metrics.